OpenStudy (anonymous):

On August 1, 2005, the national debt of the U.S. was about $7,870,000,000,000. The population of the U.S. at that time was about 297,000,000. If the national debt was divided evenly among everyone in the country, how much would each person owe? Write your answer in scientific notation.

OpenStudy (diyadiya):

\[\frac{7.87 \times 10^{12}}{2.97 \times10^{8}}\]

OpenStudy (diyadiya):

\[\frac{7.87}{2.97} \times 10 ^{4}\]

OpenStudy (anonymous):

Thanks! :D ...are you good with polynomial functions?

OpenStudy (diyadiya):

you can post the question , if i know i'll definitely answer :D

OpenStudy (anonymous):

f(x)= 5x + 2 I need to know if its a polynomial function, and write in standard form if it is.